OpenStudy (anonymous):

medal and honor Five added to 2 times a number is equal to 7 less than 4 times the number. What is the number? (Points : 1) 6 3 –6 –3

OpenStudy (anonymous):

6

OpenStudy (anonymous):

5 + 2x = 7-4x solve for x -2 = -6x x = 3

OpenStudy (anonymous):

it is not 3

OpenStudy (anonymous):

7 less than 4x is 4x-7

OpenStudy (anonymous):

Yes you are correct... 5 + 2x = 4x-7 12 = 2x x = 6 Which is why we suggest writing the solution out versus just providing answer. Thanks for correcting